H L W  F  S  U M  May 14-7:40 am May 14-9:15 am May 14-3:25 pm May 15-8:25 am May 15-9:50 pm May 16-4:50 am May 25-12:55 pmMay 25-1:45 pm May 29-9:00 pm May 29-9:15 pm June 11-6:35 am June 11-8:30 am June 11-9:45 am June 11-1:40 pm June 11-6:00 pm June 11-7:37 pm  May 21 or 22 May 24 or 25 by car Travel schedule Legend L = Lubbock H = Houston F = Frankfurt S = Samara U = Ulyanovsk M = Moscow W = Warsaw

Figure 1: This figure name...